### v1

# Toward Personalized LLM-Powered Agents: Foundations, Evaluation, and Future Directions

Abstract.
Large language models have enabled agents that reason, plan, and interact with tools and environments to accomplish complex tasks. As these agents operate over extended interaction horizons, their effectiveness increasingly depends on adapting behavior to individual users and maintaining continuity across time, giving rise to personalized LLM-powered agents. In such long-term, user-dependent settings, personalization permeates the entire decision pipeline rather than remaining confined to surface-level generation. This survey provides a capability-oriented review of personalized LLM-powered agents. We organize the literature around four interdependent components: profile modeling, memory, planning, and action execution. Using this taxonomy, we synthesize representative methods and analyze how user signals are represented, propagated, and utilized, highlighting cross-component interactions and recurring design trade-offs. We further examine evaluation metrics and benchmarks tailored to personalized agents, summarize application scenarios spanning general assistance to specialized domains, and outline future directions for research and deployment. By offering a structured framework for understanding and designing personalized LLM-powered agents, this survey charts a roadmap toward more user-aligned, adaptive, robust, and deployable agentic systems, accelerating progress from prototype personalization to scalable real-world assistants.
Personalization, Personalized LLM-powered Agent
$\dagger$} Corresponding Author.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: CSURccs: General and reference Surveys and overviewsccs: Human-centered computingccs: Computing methodologies Intelligent agents
1. Introduction
Large language models (LLMs) have evolved from passive text generators into general-purpose reasoning systems capable of understanding diverse data, planning actions, and using external tools. Building on these capabilities, LLM-powered agents extend LLMs into integrated systems that decompose complex objectives, invoke tools, interact with dynamic environments, and coordinate with humans or other agents to achieve high-level goals over extended horizons (Acharya et al., 2025; Luo et al., 2025a; Wang et al., 2024c; Tran et al., 2025). This shift marks not only a technical expansion of LLM capabilities but also a conceptual transition toward autonomous, adaptive, and socially grounded intelligent assistants.
As agent systems increasingly support long-term interaction and autonomous decision-making, personalization emerges as a central requirement for maintaining contextual continuity and aligning agent behavior with individual user goals across domains such as education, healthcare, and recommendation (Chu et al., 2025; Wang et al., 2025d; Peng et al., 2025). In LLM-powered agents, personalization extends across the full decision pipeline, shaping how agents interpret user intent, retain information over time, generate plans, and execute actions (Zhang et al., 2024e; Tseng et al., 2024). This system-wide scope introduces new challenges, including handling dynamic and multimodal feedback, maintaining consistency across interacting components, balancing user-specific adaptation with general capabilities, and safeguarding privacy and security. Correspondingly, evaluating personalized agents requires moving beyond static correctness to capture long-term effectiveness, adaptability, and user satisfaction.
Despite growing interest, reviews on personalized LLM-powered agents remain fragmented. Existing studies often concentrate on individual capabilities such as user modeling and memory construction (Wu et al., 2025c; Zhang et al., 2025i), planning and reasoning strategies (Wei et al., 2025), or adaptive interaction mechanisms (Li, 2025; Gao et al., 2025). A unified framework that connects personalization objectives across the full agent lifecycle remains limited. As a result, the interactions among different personalization mechanisms, their system-level trade-offs, and their implications for evaluation and deployment are not yet well understood.
This survey presents a capability-oriented and system-level perspective on personalized LLM-powered agents. Personalization is examined as a distributed property instantiated across four interdependent components: profile modeling, memory management, planning, and action execution. This decomposition enables a systematic analysis of how user-specific signals are represented, propagated, and operationalized throughout the agent lifecycle, from intent understanding to real-world outcomes. As illustrated in Figure 1, when an individual user submits a request, a personalized agent coordinates four interdependent capabilities to generate a tailored response. Profile representations and role definitions shape internal preference modeling, memory schemes organize and retrieve relevant contextual information, planning determines personalized decision strategies, and action execution grounds these decisions in the environment. The resulting interaction outcomes provide feedback that refines internal user representations, forming a closed loop that enables iterative and sustained personalization over time. Building on this taxonomy, we review representative methods, benchmarks, and evaluation protocols, summarize major application domains, and identify open challenges and future research directions. The survey aims to clarify the design space of personalized agents, connect benchmark-driven research with real-world deployment requirements, and provide a structured foundation for developing trustworthy, effective, and scalable personalized agent systems.
Figure 1. Overview of personalized LLM-powered agents. Upon receiving a user request, the agent coordinates profile modeling, memory, planning, and action execution to generate a tailored response. Interaction outcomes provide feedback that refines user preference representations, enabling iterative and long-term personalization.
Our contributions are as follows:
  1. (1)
    We introduce a unified taxonomy that organizes personalized LLM-powered agents around four core capabilities (profile modeling, memory management, planning, and action execution), providing a system-level view of how personalization is realized across the agent pipeline (§2).
  2. (2)
    We present an extensive review of recent techniques, benchmarks, and evaluation protocols (§3–§7), highlighting how personalization mechanisms operate within and across agent components and how they are assessed in practice.
  3. (3)
    We survey representative application domains of personalized agents and discuss open challenges and promising research directions (§89).
2. Foundations of Personalized LLM-powered Agents
2.1. LLM-powered Agents
An LLM-powered agent is an autonomous system that integrates a large language model with external tools and utilities to support step-by-step interaction with open-ended environments for task completion (Ferrag et al., 2025). The environment includes both the user and the execution context, and is characterized by an external state space $\mathcal{S}$ that governs environment dynamics and feedback.
At each time step $t$, the environment is in a state $s_{t}$ that captures the externally available information relevant to the task, including user inputs, tool outputs, or other external signals. The agent maintains an internal state $h_{t}$, which captures accumulated context such as retrieved memory, intermediate reasoning results, and internal representations used for decision making. The agent’s behavior can be abstracted as a policy that selects an action $a_{t}$ conditioned on both the current environment state and its internal state:
$\pi(a_{t}|s_{t},h_{t}).$
Actions may correspond to natural language responses, planning steps, tool invocations, or other environment interactions. The environment then evolves according to a controlled transition process, producing a subsequent external state $s_{t+1}$, and the agent updates its internal state accordingly as
$h_{t+1}=f(h_{t},a_{t},s_{t+1}).$
Starting from an initial state $s_{0}$, such as a user query or task specification, execution proceeds iteratively until a termination condition is met, yielding a trajectory
$\tau=(s_{0},h_{0},a_{0},s_{1},\ldots,a_{T-1},s_{T}).$
This trajectory reflects the agent’s evolving interpretation of the task and its interaction history. For analysis or evaluation purposes, execution outcomes can be assessed using a feedback or reward function defined over the trajectory, denoted as $r=\mathcal{R}(\tau)$.
At the system level, LLM-powered agents rely on a set of tightly coupled cognitive capabilities that enable adaptive and goal-directed behavior (Luo et al., 2025a; Wang et al., 2024c). Profiling shapes how the agent interprets its role and operating context. Memory supports continuity by retaining and organizing information across interactions. Planning governs the transformation of high-level objectives into structured decision processes. Action execution realizes these decisions through tool use or direct interaction with the environment. Together, these capabilities elevate large language models into interactive decision-making systems capable of sustained autonomy.
2.2. Personalized LLM-powered Agents
A personalized LLM-powered agent is an LLM-based agent whose internal decision-making is adapted to individual users through user-specific preferences. For each user $u\in\mathcal{U}$, the agent maintains a representation of preferences $p_{u}$, which may encode goals, constraints, habits, or behavioral tendencies inferred from interaction history, feedback, or explicit input.
At each time step $t$, the environment is in a state $s_{t}$, and the agent maintains an internal state $h_{t}^{(u)}$ specific to the user. The agent interprets and acts upon user preferences by conditioning its decision policy on $p_{u}$. This yields a user-conditioned policy
$\pi(a_{t}|s_{t},h_{t},p_{u}),$
which selects an action $a_{t}$ given the current environment state, internal state, and user preferences. The environment transitions to a new state $s_{t+1}$, and the agent updates its internal state as
$h_{t+1}=f(h_{t},a_{t},s_{t+1},p_{u}),$
where user preferences may be explicitly stored within the internal state or provided as an external conditioning signal. Executing this policy from an initial state produces a user-conditioned interaction trajectory
$\tau^{(u)}=(s_{0},h_{0}^{(u)},a_{0}^{(u)},s_{1},\ldots,s_{T}),$
where the superscript $(u)$ indicates that the trajectory is induced by decision-making conditioned on user preferences.
Personalization refers to the process through which user preferences are accumulated, represented, and integrated into an agent’s decision pipeline, allowing the same task specification $\mathcal{Q}$ to give rise to different outcomes for different users. Through repeated interaction, the agent collects user-centric data $\mathcal{D}_{u}$ and refines an internal preference representation $\hat{p}_{u}=p(\mathcal{D}_{u})$, which conditions subsequent action selection. At an abstract level, personalization can be viewed as favoring decisions that yield higher user-aligned utility over interaction trajectories:
$\pi_{u}^{*}\;\propto\;\arg\max_{\pi}\mathbb{E}_{\tau^{(u)}\sim\pi(\cdot\mid\mathcal{Q},\hat{p}_{u})}\mathcal{R}_{u}(\tau^{(u)}),$
where $\hat{p}_{u}$ denotes an internal representation of user preferences maintained by the agent, and $\mathcal{R}_{u}(\tau^{(u)})$ denotes feedback reflecting user-specific satisfaction or alignment.
From a system perspective, personalization forms a closed interaction loop. User-specific preference representations condition the agent’s perception, reasoning, and action selection; executed actions shape subsequent interactions; and the resulting outcomes provide new signals that refine $\hat{p}_{u}$. Through repeated execution of this loop, personalized LLM-powered agents progressively align their behavior with individual users, enabling long-term adaptivity and consistent user-specific behavior.
2.3. User-Centric Data
The internal representation of user preferences in LLM-powered agents is grounded in user-centric data, which captures contextual information, feedback, and interaction traces associated with a specific user. From a temporal and functional perspective, user-centric data can be broadly divided into historical data and interaction data.
Historical Data refers to information available before task execution, including user identifiers, demographic attributes, and records of prior interactions and feedback. This context provides a front-loaded prior over the user for the current episode, capturing relatively stable user characteristics together with accumulated evidence from past interactions, which supports consistent behavior and cross-task generalization.
Interaction Data is generated during task execution and captures real-time user inputs, feedback, and contextual cues. It includes both explicit signals, such as corrections or ratings, and implicit signals, such as behavioral patterns or response timing. Interaction data reflects the user’s momentary intent and is critical for guiding immediate decision-making, including reasoning adjustments and dynamic action selection.
Figure 2. User-specific data in personalization process.
These two data types enable dual-timescale personalization: historical data defines who the user is, while interaction data determines how the agent should act now. As illustrated in Figure 2, historical data guides the agent’s behavior during interaction, while interaction data refines and expands the historical record, together forming a continuous cycle of user-aware adaptation.
2.4. User Preferences
User preferences constitute the core signals that enable personalization in LLM-powered agents by conditioning their internal state, decision-making processes, and user-specific feedback. Prior work commonly categorizes preferences according to their mode of expression (Zhang et al., 2024e; Li et al., 2025c). Explicit preferences are directly specified by users and can be incorporated into the agent without additional inference. Implicit preferences are inferred indirectly from behavioral patterns or contextual cues, reflecting user judgments in a latent or under-specified form.
Beyond expression form, preferences can be characterized by their semantic function, capturing which aspect of the user they encode. We distinguish behavioral preferences, which govern how users communicate and reason, and topical preferences, which specify what users prioritize in a given context. Behavioral preferences encompass tone, reasoning style, general interaction tendencies, and personality-related traits, and are often stable across tasks. Topical preferences cover domain interests, factual stances, and likes or dislikes toward specific entities or events, and vary more with context. Both types can be stated explicitly or inferred implicitly, yielding the two-dimensional taxonomy in Figure 3.
Figure 3. Example of a two-dimensional taxonomy of user preferences. Preferences are categorized by their expression form (explicit vs. implicit) and semantic function (behavioral vs. topical), illustrated through a multi-turn conversational recommendation scenario.
In practical systems, explicit preferences can be represented in various forms, including reward vectors (Barreto et al., 2025), preference embeddings (Poddar et al., 2024), or natural-language prompts (Kim et al., 2025b). Although explicit signals are often highly informative, they are typically sparse, which limits their direct applicability for optimization. As a result, implicit preferences are frequently leveraged through in-context prompting (Kim and Yang, 2024), retrieval-augmented generation (RAG; (Qian et al., 2025b)), or preference modeling techniques (Guan et al., 2025) that extract structured representations from user behavioral feedback. When performing such an extraction, accounting for the semantic function of preferences enables more targeted modeling strategies and improves interpretability.
2.5. Capability-oriented Taxonomy
Personalization in LLM-powered agents emerges from the coordinated operation of multiple internal capabilities rather than from a single adaptation module. Given the diversity of user data, task settings, and preference types, user-specific signals must be represented, retained, reasoned over, and operationalized throughout the whole pipeline. We therefore adopt a capability-oriented taxonomy (illustrated in Figure 4) that organizes personalization into four interdependent components:
  • (1)
    Profile Modeling structures user-specific information into internal representations, defining both user characteristics and the agent’s role.
  • (2)
    Memory maintains and retrieves user-relevant information across interactions, supporting temporal continuity and consistent preference grounding.
  • (3)
    Planning integrates user representations into reasoning processes, shaping decision paths, strategy selection, and prioritization under contextual constraints.
  • (4)
    Action Execution operationalizes personalized decisions through tool invocation and environment interaction, enabling the final outcome.
{forest}
Figure 4. Taxonomy of personalized LLM-powered agents.
3. Profile Modeling
Profile modeling defines the fundamental identity of an agent, including its intrinsic attributes, behavioral tendencies, and operational boundaries (Luo et al., 2025a). In general-purpose agents, this identity is either static or pre-defined externally,established through fixed role instructions and domain expertise (Zhang et al., 2024a; Wu et al., 2024b). In personalized LLM-powered agents, however, profile modeling becomes user-aware and adaptive, incorporating information about the user and dynamically adjusting the agent’s role. To capture this bidirectional relationship, we discuss personalization from two perspectives: user profile modeling and agent role definition.
3.1. User Profile Modeling
User profile modeling extracts and organizes user-specific signals into a representation that can condition downstream personalization modules. In contrast to generic semantic context, profiles primarily encode personalization cues such as stable traits, long-term preferences, and interaction patterns that determine how an agent should respond or act for a particular user (Wu et al., 2024a). We view a user profile as a structured state constructed from heterogeneous features, where the key design questions are which signals are decision-critical and how they should be aggregated into a usable form. Accordingly, we divide profile modeling into persona-based and response-based approaches, reflecting whether the profile is anchored in user attributes or in user-specific feedback on model outputs.
Persona-based Modeling
Persona-based methods infer high-level user attributes that remain informative across contexts. Profiles may be built from explicit self-descriptions and from implicit cues in historical interactions, and are often represented as either a structured attribute vector or a free-form natural-language persona summary. For instance, AlignXpert (Li et al., 2025c) constructs a high-dimensional preference space grounded in psychological and alignment-motivated dimensions, then estimates user-specific positions to form a preference vector. FSPO (Singh et al., 2025) adopts a user-description chain-of-thought mechanism to allocate additional inference effort to preference analysis, producing more fine-grained persona descriptions. Beyond one-shot estimation, several systems iteratively refine latent user attributes through ongoing dialogue. RLPA (Zhao et al., 2025b) and Wang and Yang (2025) update user representations over interaction rounds, while RGMem (Tian et al., 2025) builds evolving profiles by extracting user insights from episodic memory fragments and progressively aggregating them via hierarchical refinement. Relatedly, RLPF (Wu et al., 2025a) optimizes concise, human-readable user summaries using a feedback loop tied to downstream performance, and CURIO (Wan et al., 2025) introduces an intrinsic motivation objective that encourages active inference of latent user types during multi-turn interaction.
Response-based Modeling
Response-based methods model user preferences through how a user evaluates candidate outputs, capturing context-sensitive notions of what constitutes a preferred response. These profiles are typically derived from fine-grained supervision such as pairwise comparisons or implicit behavioral signals, and are designed to generalize preference judgments across prompts and situations. A common approach factorizes individual preference into a shared reward feature basis with user-specific weights, enabling rapid adaptation under sparse user data. For example, RFM (Barreto et al., 2025), PReF (Shenfeld et al., 2025), and LoRe (Bose et al., 2025) represent preferences as linear combinations of shared reward features and infer user-specific coefficients from limited feedback. Complementarily, some methods parameterize user preferences in natural language to improve interpretability and controllability. Janus (Lee et al., 2024) organizes preference dimensions hierarchically from coarse categories to finer-grained value descriptions, while PLUS (Nam et al., 2025) learns text-based preference summaries that condition downstream reward modeling for personalized scoring and generation.
3.2. Agent Role Definition
Agent role definition determines who the agent is in relation to the user, which can be divided into user-simulated agent definition and adaptive agent definition based on whether the agent is on behalf of the user or not.
User-simulated Agent Definition
In some applications, especially delegation-oriented tasks such as user or dialogue simulation, agents are instantiated to act as simulated users with pre-defined attributes and preferences. These profiles usually function as static role configurations, parameterizing agent behavior for testing or evaluation and can substantially influence task planning and execution (Newsham and Prince, 2025). This one-way profile modeling emulates user traits without learning from real interaction, but remains crucial for creating realistic human–agent scenarios, enabling efficient large-scale experimentation, and ensuring behavioral diversity across synthetic users(Wang et al., 2025b; Li et al., 2025b; Gao et al., 2024b).
Adaptive Agent Definition
In interactive personalized systems, the agent role itself becomes conditional on the user profile. The agent dynamically adjusts its persona, tone, or autonomy based on the user’s characteristics, goals, and trust level, enabling a two-way adaptive process. LD-Agent (Li et al., 2024a) adopts a bidirectional user-agent modeling approach (Xu et al., 2022), utilizing a tunable persona extractor to manage long-term persona bank for both sides. PersonaAgent (Zhang et al., 2025c) optimizes the user-specific system prompt through textual loss feedback between simulated and ground-truth responses, enabling test-time user-preference alignment. PersonaX (Shi et al., 2025c) performs offline clustering and multi-persona profiling, combining with profile retrieval for online decision making. RPP (Mao et al., 2025) proposes an instance-wise prompt optimization framework, which includes both the user historical information and the task guidance for agent, aiming to generate personalized discrete prompts for better individual recommendation performance.
3.3. Discussion
Role of Profile Modeling
Profile modeling provides the foundational layer of personalization by defining both who the user is through user profiling and who the agent becomes through agent role definition. These two components offer complementary but distinct perspectives: user profiling seeks to construct expressive representations of user traits and preferences, while agent role definition determines how the agent’s operational style should adapt to those user characteristics.
Challenges
Despite their tight conceptual coupling, they face different technical challenges. On the one hand, user profiling is challenged by sparse, noisy, and inconsistent preference signals, as well as the potential conflict between long-term tendencies and context-specific preferences. On the other hand, agent role definition faces the trade-off between flexibility and stability across tasks, avoiding both rigid roles and undesirable role drift. Finally, a broader challenge lies in the weak coupling between profiling and role adaptation. Effective personalization requires bidirectional alignment where the agent must update its role based on refined user profiles, while user profiling should leverage signals arising from the agent’s adaptive behavior.
Future Directions
Existing user profiling methods often yield representations that are static or task-specific, and agent roles that lack controllability or principled adaptation dynamics. Moreover, user profiling and agent role definition are typically developed in isolation, with limited mechanisms for mutual adjustment. Going forward, promising directions include lightweight and interpretable profiling methods that remain robust under sparse signals, more unified frameworks that couple evolving user profiles with stable yet adaptive agent roles.
4. Memory
While profile modeling captures stable and high-level user traits, many fine-grained details essential for personalized responses cannot be fully encoded in a plain profile. To support such detailed personalization, agents rely on a memory module, which refers to the system’s ability to retain, recall, and leverage past interactions to improve future behavior (Weng, 2023). Memory is commonly divided into personal memory, which stores user inputs and interaction histories, and system memory, which records intermediate reasoning states or task-execution artifacts (Wu et al., 2025c). This survey focuses on personal memory, as it directly shapes how agents adapt to individual users. In the following, we first discuss how personal memory is structured and updated, and then how it is retrieved to support personalized behaviors.
Personal memory can be implemented as internal memory, which embeds user-related information within the LLM’s internal components, or as external memory, which stores information in an auxiliary system. Internal memory approaches operate through modifications to model parameters (Ning et al., 2025a; Wang et al., 2025c; Jin et al., 2025a; Tan et al., 2024), key–value caches (Xi et al., 2024; Qian et al., 2025b), or hidden states (Zhang et al., 2025e, j; Wang et al., 2025e). Although parameter-efficient and compact (Wang et al., 2024e), internal memory is constrained by limited memory capacity, challenges in updating stored information, and reliance on model retraining, making it less suitable for dynamic, long-horizon personalization. By contrast, external memory stores information in an auxiliary system, usually combined with retrieval-augmented generation (RAG; Lewis et al., 2020), which enriches the input prompt with information retrieved from the external knowledge base, thereby improving user satisfaction and engagement. Compared to internal memory, external memory is more flexible and especially beneficial for users with limited personal data (Salemi and Zamani, 2025a). However, it requires careful design for storage, management, and personalized access. Given these considerations, we focus on external personal memory systems and structure this section around two key components: (1) personal memory schemes, which define how user-related information is stored and updated, and (2) personalized retrieval, which determines how stored information is accessed to guide personalized behaviors.
4.1. Personal Memory Schemes
Personal memory schemes define the foundation of the memory module, specifying how user-related information is organized, stored, and updated throughout interaction. Based on the timescale and influence of stored information, personal memory is often divided into short-term memory, which captures recent conversational context, and long-term memory, which accumulates enduring user information across sessions. Short-term memory provides reliable signals for immediate responses but is limited by context window constraints and cannot effectively utilize information beyond the current interaction session. On the other hand, long-term memory enables representation of long-tail individual data and supports self-evolution through accumulated experience (Jiang et al., 2024b), but suffers from information overload and difficulties in timely updating. As a result, recent personalized agents increasingly adopt hybrid memory designs that combine both short- and long-term memory to balance continuity and fine-grained adaptation (Li et al., 2024a), motivating careful design of the underlying memory structure and update mechanism.
4.1.1. Memory Structure
The memory structure determines what information is stored and how it is represented, forming the foundation for effective personalized retrieval. Early agent systems often appended full dialogue histories or execution trajectories to memory (Liu et al., 2024b), but such raw storage introduced redundancy, limiting the agent’s ability to surface user-relevant information. To achieve scalable personalized memory management, recent works convert interaction histories into compact and structured representations tailored to task demands (Liu et al., 2023; Zhong et al., 2024). Based on the format, memory structures can be broadly categorized as textual memory and structured memory.
Textual Memory
Textual memory stores user-related information in natural language form, typically by summarizing multi-turn interactions into concise and interpretable units. Because such representations preserve rich semantics and align naturally with the input–output format of LLMs, textual memory has become the dominant choice for personalized agents across diverse tasks (Zhang et al., 2025i). However, effective textual memory design faces two key challenges: defining appropriate segmentation boundaries and capturing comprehensive user-related information. Early approaches often relied on rigid turn-level or session-level segmentation, which frequently misaligned with the semantic units of dialogue and resulted in fragmented retrieval. To address this, recent work has explored more flexible, topic-coherent segmentation strategies. For instance, SeCom (Pan et al., 2025) and RMM (Tan et al., 2025c) partition conversations into semantically coherent segments, enabling memory units to reflect stable conversational topics rather than arbitrary boundaries. Nemori (Nan et al., 2025) further improves segmentation accuracy by using a boundary detector to group entire interactions into coherent episodes, preserving the integrity of user intent. Beyond segmentation, many systems enrich textual memory with auxiliary attributes such as timestamps, topic labels, and user personality (Liu et al., 2023; Xu et al., 2025d; memodb, 2025). For example, CAIM (Westhäußer et al., 2025a) introduces a tagging system that stores events with tuples of related tags, inductive thoughts, and timestamps. MemoryBank (Zhong et al., 2024) and PersonaAgent (Zhang et al., 2025c) adopt hierarchical summarization to model multi-level event structure and dynamic personality shifts, simulating the multi-component structure of human memory and enabling more coherent and human-like personalized interaction. MemInsight (Salama et al., 2025) is an autonomous memory augmentation approach that employs an attribute mining module to extract structured, semantically meaningful attributes from input dialogues. Several recent works draw inspiration from human memory systems (Shan et al., 2025; Zhang et al., 2025f). These systems typically distinguish between semantic memory and episodic memory (Liu et al., 2025c; Pink et al., 2025; Long et al., 2025). MIRIX (Wang and Chen, 2025) proposes an architecture with six specialized memory components (core, episodic, semantic, procedural, resource, and knowledge vault) organized hierarchically. Mem-$\alpha$ (Wang et al., 2025f) proposes an architecture with core, semantic, and episodic memory modules, combined with an RL framework that teaches agents to manage complex memory through interaction and feedback. MemWeaver (Yu et al., 2025) combines temporal and semantic dimensions across behavioral-level and cognitive-level memory to jointly represent user actions and long-term preferences. Collectively, these techniques advance textual memory toward richer representations, more coherent segmentation, and multi-level organization, allowing personalized agents to retain user-relevant details more faithfully and capture both explicit and implicit signals with higher accuracy.
Structured Memory
Structured memory provides a more explicit and computationally manipulable representation of interaction histories by storing information in predefined formats enriched with structural relationships. Compared with textual summaries, structured memory introduces stronger inductive biases that support scalable retrieval, multi-granularity reasoning, and more fine-grained personalization. Current approaches can be broadly grouped into vector-based memory, hierarchical tree structures, and graph-based memory architectures. Vector-based memory represents each memory unit using learned embeddings stored in a vector database, enabling scalable similarity search and fast retrieval of user-relevant information. This design has become a standard practice in personalized agents due to its efficiency and flexibility (Ocker et al., 2025; Chhikara et al., 2025; Wu et al., 2025d). However, vector memory alone discards structural relationships between memory items and offers limited interpretability, motivating more structured alternatives or hybrid systems that combine vectors with explicit organizational formats. Hierarchical tree structures organize memory at multiple abstraction levels, where each node corresponds to a memory unit and parent–child relationships encode increasingly coarse summaries. For instance, RAPTOR (Sarthi et al., 2024) recursively clusters and summarizes text into a tree, supporting retrieval at both fine- and coarse-grained levels. MemTree (Rezazadeh et al., 2024) similarly constructs a memory hierarchy in which each node stores aggregated textual content and semantic embeddings, allowing memory to be represented at varying abstraction depths. Graph-based architectures capture richer relationships by linking memory units through typed edges representing temporal, semantic, or relational dependencies (Jimenez Gutierrez et al., 2024; Li et al., 2025i; Chhikara et al., 2025). Some methods focus on behavioral coherence, such as MemWeaver (Yu et al., 2025), which builds an event-level behavior graph where interaction events are nodes connected by temporal and semantic edges, along with a context-aware random walk mechanism. Other designs integrate multi-dimensional subgraphs to capture more nuanced relationships between memory units. For example, AriGraph (Anokhin et al., 2024) combines semantic knowledge (entity–relation triplets) with episodic observation nodes, enabling both long-term knowledge accumulation and temporally grounded recall. Zep (Rasmussen et al., 2025) introduces a temporally-aware dynamic knowledge graph with three tiers, including episodic for non-lossy data storage, entity-level semantic for extracted entities and relationships, and community-level conceptual for high-level summarizations of strongly connected entity clusters. MemGAS (Xu et al., 2025a) further constructs cross-granularity associations across session-, turn-, summary-, and keyword-level memory to enrich personalized retrieval signals. While powerful, these graph-based structures often depend on computationally expensive LLM-based entity and relation extraction and may lose fine-grained contextual information during graph construction. To address this, SGMem (Wu et al., 2025d) proposes a lightweight sentence-graph memory, linking chunk nodes with their constituent sentences and connecting sentences via similarity edges to better preserve contextual semantics. Similarly, LiCoMemory (Huang et al., 2025g) introduces CogniGraph, a lightweight hierarchical indexing graph that organizes session summaries, entity–relation triples, and dialogue chunks through cross-layer hyperlinks, keeping the graph minimal while storing full content externally. Across these approaches, structured memory aims to provide richer organization by modeling how information relates, evolves, and aggregates across interactions.
4.1.2. Update mechanism
Memory update mechanism determines how an agent’s memory evolves as new interactions occur, ensuring that stored information remains accurate, consistent, and aligned with the user’s changing preferences. When a new memory unit arrives, the system must decide whether to incorporate it as a new entry, revise or merge existing ones, or prune outdated or contradictory content (Fang et al., 2025). Broadly, memory updating differs across two types of external memory: textual memory, where updates rewrite or reorganize summary-level information, and structured memory, where updates modify entity–relation structures through node/edge creation, conflict resolution, and temporal maintenance. We review the update mechanism for these two categories separately.
Textual memory stores interaction histories as natural-language segments or summaries. Updating therefore concerns how these segments are merged, rewritten, or reorganized when new dialogue arrives. Existing approaches can be grouped into similarity-driven updates and inference-guided updates.
Similarity-driven updates
These methods update memory by comparing new segments with existing summaries and adjusting them based on semantic proximity and auxiliary metadata such as timestamps or topic labels. RMM (Tan et al., 2025c) summarizes each session by topic and updates the memory bank through semantic merging, replacing or consolidating summaries with high topical similarity. CAIM (Westhäußer et al., 2025a) extends this paradigm with a post-thinking step that generates structured, tag-annotated summaries from recent dialogue and integrates them into memory while refining related historical entries. Mem0 (Chhikara et al., 2025) introduces an LLM-driven update framework that chooses among ADD, UPDATE, DELETE, and NOOP operations for each new fact based on semantic similarity with existing memories. Memory-R1 (Yan et al., 2025a) builds on this with a reinforcement-learning memory manager that learns the optimal operation policy across multi-session interactions.
Inference-guided updates
Other systems treat memory updating as a reasoning or decision-making process, allowing the agent to incorporate implicit preference shifts or contextual dynamics not captured by similarity alone. Nemori (Nan et al., 2025) exemplifies this through its Predict–Calibrate mechanism, achieving proactively evolving memory based on gaps between predicted and actual user responses. PREMem (Kim et al., 2025c) similarly performs pre-storage reasoning across sessions to analyze thematic evolution before inserting new content, enabling more consistent long-term topic tracking. These methods provide more adaptive and self-correcting memory dynamics, capturing both short-term changes and long-term shifts in user preferences.
Structured memory organizes information into interconnected architectures where inserting a new unit often requires reorganizing its related counterparts. Update mechanisms can likewise be grouped into similarity-driven and reasoning- or agentic-driven strategies.
Similarity-driven updates
Most personalized structured-memory systems update the memory by assessing semantic similarity between new information and existing units, then making local modifications to maintain coherence. In hierarchical tree structures, MemTree (Rezazadeh et al., 2024) updates memory by traversing the tree via semantic similarity, inserting new nodes through sibling attachment or leaf expansion, and incrementally updating ancestor nodes with LLM-based aggregation to preserve hierarchical coherence. Graph-based architectures demand even more complex updates, as every node may connect to multiple others (Huang et al., 2025g). $\text{Mem0}^{g}$ (Chhikara et al., 2025) embeds new triples, matches them to existing nodes, and establishes metadata-rich edges while using an LLM-based conflict resolver to deactivate outdated relations. Similarly, Zep (Rasmussen et al., 2025) integrates new episodic facts by timestamp alignment, updating or deactivating inconsistent edges and adjusting higher-level community clusters to maintain a temporally grounded semantic–episodic graph. MemGAS (Xu et al., 2025a) clusters new and historical multi-granularity chunks with GMM, selectively reinforcing only contextually relevant associations to ensure stable personalization over long horizons.
Reasoning- or agentic-driven updates
A smaller but emerging set of systems incorporate explicit reasoning, decision-making, or agentic inference to guide how memory should be updated, moving beyond similarity thresholds toward more cognitively aligned update dynamics. A-mem (Xu et al., 2025d) adopts an agentic update scheme in which each new memory becomes a structured note that triggers semantic linking and retroactive refinement of historical notes, enabling a continuously self-evolving memory graph. Although not designed for personalization, recent non-personalized graph-memory systems illustrate the potential of this direction. TRAIL (Zhao et al., 2025c) performs inference-driven updates by validating and refining candidate triples through confidence-guided decisions rather than pure similarity matching, while AriGraph (Anokhin et al., 2024) uses agentic reasoning to determine whether new observations should create links, update existing nodes, or refine episodic associations. These systems demonstrate that reasoning-based graph evolution can produce more coherent and cognitively grounded memory updates, suggesting a promising direction.
4.2. Personalized Retrieval
Memory retrieval determines which stored information becomes available to an agent at inference time and therefore directly influences the quality and consistency of personalized behaviors. Unlike conventional RAG systems, retrieval for personalized LLM-powered agents must simultaneously satisfy three requirements: (1) content relevance to the current query, (2) structural consistency with the underlying memory organization, and (3) personal alignment, ensuring that retrieved information reflects the user’s preferences. Existing methods address these challenges through three dominant mechanisms: content-based retrieval, structure-aware retrieval, and policy-guided retrieval.
Content-based retrieval
Content-based retrieval identifies relevant memory by measuring semantic or lexical similarity between the user’s query and stored information, typically implemented with dense retrievers (e.g., FAISS (Johnson et al., 2019), Contriever (Izacard et al., 2021)) or sparse lexical retrievers (e.g., BM25 (Robertson et al., 2009)). Retrieved segments are generally appended to the agent’s prompt to enrich contextual grounding (Wang and Chen, 2025; Kim et al., 2025c; Xu et al., 2025d). To capture comprehensive information about the user, content-based retrieval can be operated at multiple granularity levels. For example, RAPTOR (Sarthi et al., 2024) and MemTree (Rezazadeh et al., 2024) perform hierarchical content-based retrieval by selecting nodes whose embeddings are most similar to the query across different abstraction levels, even though the underlying tree structure is not explicitly traversed for reasoning. To further improve retrieval quality, some methods incorporate post-retrieval filtering or query refinement. Yuan et al. (2025) adopt a self-reflection mechanism that assesses whether the retrieved content is sufficient for response generation and refines the query when necessary. While effective at capturing semantic relevance, content-based approaches may overlook structural dependencies and long-range contextual relationships, motivating more structured retrieval designs.
Structure-aware retrieval
Structure-aware retrieval leverages the organization of memory, such as graph connectivity and episodic relationships, to retrieve information that is contextually linked to the query rather than merely similar in content. Such methods are often implemented in hierarchical or hybrid modes, where an initial retrieval of entities, topics, or relations guides subsequent retrieval of detailed memory units (Jimenez Gutierrez et al., 2024; Gutiérrez et al., 2025), potentially augmented with graph-based techniques such as Personalized PageRank (PPR; Bahmani et al., 2010). For instance, AriGraph (Anokhin et al., 2024) first retrieves semantically relevant relational triplets and then, conditioned on these triplets, retrieves the most relevant episodic memories. Zep (Rasmussen et al., 2025) combines semantic search with graph-structured signals, such as episode-mention frequency and node-distance scoring, to identify contextually relevant conversational episodes within its temporal memory graph. Other approaches enhance structure-aware retrieval with adaptive routing or self-reflection. MemGAS (Xu et al., 2025a) computes multi-granularity semantic similarities and integrates them using an entropy-based router, then applies PPR over a dynamically maintained association graph to retrieve structurally coherent memory units. AssoMem (Zhang et al., 2025b) performs associative retrieval by selecting clue nodes and linked utterances, followed by a multi-signal ranking procedure that fuses relevance, graph-based importance, and temporal alignment through a mutual-information–driven fusion strategy. By explicitly modeling the structural dependencies among memory units, structure-aware retrieval provides richer contextual grounding and improves the coverage of long-range or implicit user preferences.
Policy-guided retrieval
Policy-guided retrieval refers to retrieval strategies in which the selection of information is governed by user- or task-conditioned decision policies rather than static similarity matching (Zhong et al., 2024; Zhang et al., 2025c). Unlike content-based or structure-aware retrieval, which rely primarily on semantic proximity or memory topology, policy-guided retrieval actively determines how retrieval should proceed and can operate over arbitrary corpora without relying on a persistent user memory store. A first line of work focuses on pre-retrieval query transformation, where the retrieval intent is personalized before search. PBR (Zhang et al., 2025g) generates user-style pseudo feedback and corpus-anchored expansions to form a personalized query representation, while PersonaRAG (Zerhoudi and Granitzer, 2024) adapts document selection and ranking using real-time behavioral signals such as reading patterns and interaction context. A second line of work adopts policy-driven strategy selection, learning to invoke or integrate heterogeneous retrieval tools based on the task. UniMS-RAG (Wang et al., 2024b) formulates retrieval as a policy-learning problem, where supervised and reinforcement signals train the system to choose among keyword search, dense retrieval, and knowledge-graph lookup and to combine their results effectively. Persona-DB (Sun et al., 2025a) employs a two-stage persona-aware policy that first matches users via high-level persona keys and then combines retrievals from target users and collaborator users using a learned composition ratio to address sparsity and enhance personalization. A third direction incorporates schema-guided information completion, framing retrieval as a structured inference process. OCG-Agent (Shi et al., 2025b) decomposes narrative queries into an abstract data type schema and triggers a multi-route retrieval routine that fills schema fields in parallel; a reflect-and-complete loop then issues targeted supplementary retrievals until all required attributes are satisfied, followed by semantic normalization to unify heterogeneous results. Collectively, policy-guided retrieval shifts retrieval from passive document matching toward an active, decision-centric paradigm that adapts retrieval behavior to user preferences, task requirements, and interaction context, enabling more flexible and personalized information acquisition.
4.3. Discussion
Role of Personal Memory
Personal memory bridges short-term intent and long-term preference modeling. It fulfills three closely related roles. First, it supports preference evolution, allowing the agent to track how user preferences change across sessions rather than relying solely on fixed trait representations. Second, it provides personalized contextual grounding, preserving fine-grained user details that are essential for accurate personalization but cannot be fully encoded in static profiles. Third, it enables preference-conditioned generation, guiding the agent to interpret new queries and produce responses informed by retrieved user-relevant signals. Together, these roles transform memory from a passive log of interactions into an active component of the personalization loop.
Challenges
Despite its importance, personal memory faces several challenges. First, user preferences evolve at different rates, requiring memory systems to balance rapid adaptation with the preservation of long-term traits. Second, memory stores may accumulate redundancy or hallucinations, especially when updates rely on LLM-driven summarization or relation extraction, leading to degraded personalization over long horizons. Third, ensuring that retrieved content is not only semantically relevant but also truly aligned with user preferences remains difficult. Finally, personal memory raises privacy and transparency concerns, highlighting the need for user-controllable and privacy-preserving memory mechanisms.
Future Directions
Memory modules also have inherent limitations. Textual memories scale poorly and require repeated rewriting; structured memories provide controllability but depend on costly extraction and risk losing contextual nuance; update mechanisms oscillate between shallow similarity rules and computationally heavy inference; and retrieval pipelines remain sensitive to ranking biases and distribution shifts. These limitations point to several broad future directions across the memory pipeline, including hybrid memory architectures that combine the semantic richness of textual representations with the organizational clarity of structured formats, more principled and user-controllable update mechanisms that ensure reliability and transparency, and improved retrieval strategies that better integrate personal alignment with semantic relevance.
5. Planning in Personalized LLM-powered Agents
Planning constitutes the cognitive engine of intelligent agents, enabling the decomposition of high-level objectives into actionable steps and the formulation of execution strategies (Luo et al., 2025a; Sun et al., 2023). In the context of personalized agents, planning transcends generic task completion; it must optimize for user utility, aligning decision paths with individual preferences, constraints, and latent intents (Zhang et al., 2024e; Gao et al., 2024a; Han et al., 2025b). Unlike standard planners that prioritize efficiency or logical correctness, personalized planners navigate a constrained search space defined by the user’s unique profile and behavioral history, often requiring the integration of long-term memory and preference models (Westhäußer et al., 2025a; Huang et al., 2025c). We categorize these methodologies into two distinct paradigms based on the integration of user signals: One-shot Planning, where personalization is internalized as a prior during reasoning, and Feedback-driven Planning, where personalization emerges dynamically through interaction (Zhang et al., 2024d; Suri et al., 2025).
5.1. One-Shot Planning
One-shot personalized planning refers to generating a complete plan in a single inference pass, where a plan is represented as a structured decomposition of subgoals and decisions conditioned on user-specific constraints and latent intent. We analyze one-shot methods along two axes: how user signals are grounded into planning representations (Preference Induction and Memory Conditioning) and how plans are synthesized via compute-intensive generation (Internal Refinement).
Preference Induction.
This line of work constructs a planning-ready user state by inferring latent intent and translating heterogeneous user signals into explicit constraints or soft objectives that guide plan generation (Han et al., 2025b). For instance, ALIGNXPLORE (Li et al., 2025d) utilizes a User-description Chain-of-Thought framework to synthesize a compact, semantically grounded preference description from sparse interaction traces. Instead of retrieving raw logs, the model performs extended inductive reasoning to form global preference constraints that condition downstream planning. Similarly, in sequential recommendation, ReaRec (Tang et al., 2025) proposes a Think-Before-Recommend paradigm, using reasoning position embeddings to infer the user’s latent intent trajectory prior to plan generation. In embodied and decision-making contexts, PbP (Xu et al., 2025b) demonstrates that treating learned user preferences as intermediate abstractions substantially improves few-shot personalized planning performance.
Memory Conditioning.
Complementary to inductive constraint formation, memory-grounded methods condition one-shot planning on selectively retrieved and compressed user history, ensuring consistency with both long-term preferences and recent context (Tan et al., 2025c; Westhäußer et al., 2025a). PRIME (Zhang et al., 2025f) integrates a dual-memory architecture with a Personalized Thought Process that explicitly traces how past experiences and stable beliefs influence current decisions. Before plan generation, PRIME synthesizes a personalized thought trace via self-distillation, effectively grounding planning in the user-specific cognitive context. PersonaAgent (Zhang et al., 2025c) adopts a retrieval-augmented persona mechanism, dynamically constructing system prompts from a unified memory bank to maintain cross-session planning consistency. Related work in recommendation and dialogue systems similarly shows that structured user memory retrieval significantly improves alignment and coherence in one-shot decision-making (Huang et al., 2025c; Chen, 2025).
Internal Refinement.
Beyond direct one-pass generation, internal refinement improves it by constructing an initial plan and then internally critiquing and revising it against a user-conditioned objective. Related self-refinement templates such as Reflexion and structured critique-and-revise frameworks provide general mechanisms for generating reflective feedback and triggering plan edits (Shinn et al., 2023; Gou et al., 2023), while recent test-time-compute reasoning models make deeper within-pass self-correction increasingly practical (Muennighoff et al., 2025; Guo et al., 2025). Representative systems instantiate this paradigm in several ways. In personalized travel planning, TravelPlanner++ (Singh et al., 2024) and PTS (Shao et al., 2025) study reflection-style itinerary refinement, where intermediate plans are iteratively revised under derived user implicit long-horizon preferences. PREFINE (Ueda and Takayanagi, 2025) constructs a pseudo-user critic and user-specific rubrics from interaction history, then critiques and revises candidates to better satisfy the target user. In embodied command execution, IoTGPT (Yu et al., 2026) applies self-correction by testing intermediate command sequences in a simulated environment and revising the plan before deployment.
5.2. Feedback-Driven Planning
Feedback-driven planning treats the initial plan as a hypothesis to be refined through interaction. These systems recognize that user intent is often underspecified, evolving, or partially latent (Liu et al., 2025d; Zhang et al., 2024d), and employ iterative loops to converge on an optimal personalized solution (Suri et al., 2025; Wang et al., 2023b). We further organize feedback-driven planning by the source of feedback and how it updates the plan, including information-seeking clarification, mixed-initiative user revision, and non-user feedback signals.
Clarification Seeking
A central challenge in feedback-driven personalized planning lies in deciding when to ask for additional information: excessive clarification increases user burden, while premature commitment risks misinterpreting user-specific constraints or latent intent. Prior studies indicate that user requests in planning tasks are frequently underspecified or ambiguous (Liu et al., 2025d), motivating approaches that treat clarification as decision-making under uncertainty. For instance, SAGE-Agent (Suri et al., 2025) formulates selective questioning as a POMDP and issues clarification queries only when the expected value of information outweighs interaction cost. Similarly, ECLAIR (Murzaku et al., 2025) focuses on detecting missing arguments or ambiguous entities in user instructions and triggers targeted follow-up questions before committing to a plan. Additionally, Ask-before-Plan style frameworks (Zhang et al., 2024d) explicitly decouple clarification from plan synthesis, showing that resolving key uncertainties upfront improves robustness and downstream plan quality. Collectively, these methods establish clarification as a principled and integral component of feedback-driven personalized planning rather than an ad hoc interaction heuristic.
User Revision.
User edits provide a high-fidelity personalization signal that can update user-conditioned objectives, constraints, and preferences for subsequent planning and generation. Gao et al. (2024a) show that learning latent preferences from historical edits can reduce future edit effort and improve user-specific alignment in subjective settings. Building on this mixed-initiative view, AIPOM (Kim et al., 2025a) represents plans as editable structures, enabling users to revise intermediate plans while the agent updates constraints and priorities accordingly. Similar revision loops appear in domain systems such as Dango (Chen et al., 2025a), which supports iterative correction of agent-produced data-processing workflows, and PlanFitting (Shin et al., 2025), which refines personalized exercise plans through conversational revisions. This revision-driven interaction pattern is also adopted in deployed coding agents, where users review proposed changes and iteratively provide follow-up edits or instructions to steer subsequent updates (Anthropic, 2025; Github, 2025).
Non-user Feedback.
Beyond direct user input, plans can be refined using non-user signals that improve feasibility and preference satisfaction without repeatedly burdening the user. One line of work leverages agent-mediated feedback: VAIAGE (Liu et al., 2025a) employs specialized agents to negotiate route feasibility and recommendations under user preferences, while Collab-Rec (Banerjee et al., 2025) supports multi-stakeholder planning with explicit user-advocacy mechanisms. Another line exploits environment-mediated feedback by simulating intermediate outcomes. For instance, SceneWeaver (Yang et al., 2025) follows a Reason–Act–Reflect loop and triggers re-planning when rendered intermediate results violate user-specified attributes.
5.3. Discussion
Role of Personalized Planning.
Personalized planning serves two complementary roles in LLM-based agents. First, it operationalizes personalization by bridging upstream user intent inference and preference modeling into downstream decision-making (Chen, 2025), instantiating user-contingent objectives, constraints, and trade-offs that govern a plan. Second, particularly in feedback-driven settings (Kim et al., 2025a), planning functions as a closed-loop optimization process that alternates between eliciting informative feedback and revising the plan, enabling the agent to proactively and progressively uncover user intent and converge toward a higher degree of personalization.
Challenges.
Personalized planning raises several challenges. First, user intent and preferences are often partially observed and evolve over time (Suri et al., 2025; Liu et al., 2025d), complicating when to infer, elicit, and commit. Second, personalized utility is also multi-objective (Shao et al., 2025), requiring arbitration among competing constraints and idiosyncratic trade-offs. Moreover, interactive refinement must remain controllable and convergent, limiting user burden while avoiding unstable revisions. Existing paradigms involve inherent trade-offs. One-shot planning internalizes user signals as priors for low-latency, coherent plans, but can be brittle under sparse or shifting signals. Feedback-driven planning improves alignment via iterative elicitation and revision, at the cost of additional latency and user effort (Liu et al., 2025a).
Future Directions.
Future work may benefit from developing uncertainty-aware planning workflows that adaptively allocate internal computation and revision effort based on confidence in inferred preferences and task stakes (Muennighoff et al., 2025), combining the efficiency of one-shot planning with the alignment benefits of feedback-driven refinement. Besides, a promising direction is to learn user-conditioned planning patterns at both inter- and intra-user levels (Qiu et al., 2025a), capturing systematic differences in decomposition style and trade-off resolution while modeling their stability and drift over time, therefore providing reusable personalization signals that generalize across tasks. Finally, several aspects remain comparatively under-explored, including robustness to variations in user specification (Suri et al., 2025) and privacy-preserving planning (Wu et al., 2025e) that supports long-horizon user conditioning without exposing raw personal data.
6. Action Execution in Personalized LLM-powered Agents
Action execution is the stage where a personalized agent realizes decisions through concrete operations in the external environment and closes the loop with execution feedback (Shen, 2024; Kim et al., 2024). Unlike purely task-centric execution, this layer must ground actions in user-specific constraints and preferences, and remain adaptive when tools or environments produce unexpected outcomes. We organize personalized action execution into two stages: a Pre-action Stage that selects and parameterizes actions under user-conditioned constraints, and a Post-action Stage that leverages execution signals for recovery and preference-consistent outcome realization.
6.1. Pre-action Stage
The pre-action stage focuses on user-conditioned action decision and grounding. It translates a selected intent or plan step into executable tool calls or environment operations under user-specific constraints. We further decompose this stage into action policy, which selects among functionally valid execution options for the same step, and action grounding, which instantiates the selected action with user-specific arguments and realizations.
6.1.1. Action Policy
At this level, personalization manifests as a bias over execution choices, spanning both tool utilization and higher-level action decisions. PEToolLLaMA (Xu et al., 2025c) is an early effort that formalizes personalized tool learning from interaction history, constructs a tool-use dataset with diverse user preferences, and trains models to improve personalized tool selection. In the web-agent setting, PUMA (Cai et al., 2025) adopts a similar preference-aware learning strategy, showing that personalization can also shape action choices during web navigation and task completion. ToolSpectrum (Cheng et al., 2025) further incorporates user profiles together with environmental context, indicating that context-aware conditioning can improve user-aligned tool utilization. Beyond selecting appropriate tools, ColorAgent (Li et al., 2025f) learns to proactively engage the user to clarify ambiguous intentions or incomplete instructions. Moving beyond tool invocation, ValuePilot (Luo et al., 2025b) studies action-level decision making under user values, where agents evaluate and select actions according to user value preferences to produce context-sensitive and individualized choices.
6.1.2. Action Grounding
Action grounding operationalizes a selected action by instantiating it into executable tool calls with user-conditioned constraints and parameters (Cai et al., 2025; Cheng et al., 2025). While many requests can be grounded by directly applying user-specific settings, personalization becomes more challenging in corner cases. A frequent grounding-time failure mode is infeasibility, where the instantiated query becomes unsatisfiable under the imposed constraints. AWARE-US (Kurmaz, 2026) frames this as a preference-aware query repair problem, arguing that agents should restore feasibility by relaxing the least preferred constraints to the user rather than applying default heuristics. Another recurring issue is the omission of essential tool arguments. Huang et al. (2025d) study how agents infer unspecified arguments from user profiles, reducing execution friction while maintaining alignment with user intent.
6.2. Post-action Stage
The post-action stage closes the loop after an action is executed. Given execution feedback, the agent evaluates whether the outcome satisfies user-conditioned objectives, applies corrective adjustments when mismatches arise (execution-time correction), and produces final outputs that conform to user preferences and quality criteria (outcome realization).
6.2.1. Execution-Time Correction
A failed execution commonly produces feedback signals, which can be leveraged for adjusting behavior without re-planning from scratch. While explicit studies on preference-aware recovery in LLM-powered agents remain limited, adjacent work on mitigating tool overuse under diminishing returns (Qian et al., 2025a) and structured reflection against tool-interaction failures (Su et al., 2025) provides foundational mechanisms upon which personalized recovery strategies can be built.
6.2.2. Outcome Realization
Even when multiple execution outcomes are functionally valid, user satisfaction depends on how results are presented in accordance with individual preferences. In recommendation-oriented agents, personalized re-ranking is a core mechanism for transforming candidate results into ordered outputs that reflect user-specific utility (Pei et al., 2019). Methods such as PEAR and multi-level interaction reranking (MIR) explicitly model cross-item dependencies together with user preference features, yielding more faithful final rankings and improved utility metrics (Li et al., 2022; Xi et al., 2022). Beyond ranking, personalization also governs which execution outcomes are retained or discarded. WARPP (Mazzolenis and Zhang, 2025) emphasizes runtime personalization in workflow execution by pruning or adjusting conditional branches based on user attributes, reducing irrelevant or erroneous outcomes without re-planning the entire process. In addition to selecting among existing results, outcome realization may involve proactive augmentation. ETAPP (Hao et al., 2025) treats proactivity as a personalization signal, evaluating whether agents can propose additional actions or suggestions to better satisfy user needs.
6.3. Discussion
Role of Personalized Action Execution
Action execution constitutes a core stage that bridges high-level reasoning with tangible outcomes in LLM-based agents (Qu et al., 2025; Chowa et al., 2026). By determining how abstract decisions are instantiated, execution-level choices directly influence whether personalization is effectively realized rather than remaining at the level of internal reasoning, which substantially shapes user experience, trust, and perceived utility. Consequently, personalized action execution functions as the final bottleneck of personalization in agent systems, translating inferred user intent into observable behavior and closing the loop between internal personalization mechanisms and external outcomes.
Challenges
Designing personalized action execution mechanisms remains challenging due to several factors. First, execution contexts are highly heterogeneous across tools, environments, and domains, complicating the design of general mechanisms that consistently respect individual preferences. Second, personalization signals at execution time are often subtle, requiring abstract preferences to be translated into concrete parameters rather than explicit instructions. Third, execution feedback may expose ambiguity or brittleness in how user preferences were interpreted upstream, necessitating careful handling and, in some cases, coordination with planning or preference modeling components to avoid cascading errors.
Future Directions
Research on personalized action execution remains relatively limited, leaving several promising directions open. At the execution primitive level, developing personalization-aware tools that emphasize non-functional characteristics could enable differential behavior beyond task correctness. At the control level, learning preference-conditioned execution policies that generalize across tools and environments may allow agents to reuse personalization signals more effectively. Additionally, agents should better handle vague or missing arguments, localize execution failures, and aggregate outcomes in ways that reflect user priorities rather than default heuristics. Finally, there is a growing need for execution-level evaluation protocols that move beyond task success to assess preference adherence, consistency, and user satisfaction, in line with recent calls for more holistic evaluation of agent capabilities.
7. Evaluation
Evaluating personalized agents presents a unique set of challenges that extend beyond the traditional assessment of AI systems, as the target of optimization is not merely task correctness but long-horizon user-specific satisfaction. These challenges largely stem from the implicit, context-dependent, and evolving nature of human preferences and desires, which are often under-specified, situation-sensitive, and updated through interaction. Consequently, evaluation should prioritize how well an agent aligns with an individual user’s expectations and experience, rather than only how well it completes a task in an objective sense. This motivates the need for a comprehensive evaluation mechanism that characterizes personalization from general principles down to “in-one’s-shoes” assessments that reflect each user’s subjective utility. In this section, we provide a systematic overview of evaluation for personalized LLM-powered agents, with Figure 5 summarizing how evaluation goals, assessment paradigms, and benchmark families are interconnected within a unified framework.
Figure 5. Overview of evaluation for personalized LLM-powered agents. Evaluation is organized along three layers: (1) evaluation goals and metric dimensions, including effectiveness, adaptivity, generalization, robustness, and risk; (2) assessment paradigms, such as automatic scoring, rule-based checking, learned evaluators, and LLM-as-a-judge; and (3) representative benchmark families, including interactive alignment and user-substitution settings.
7.1. Evaluation Goals and Metrics
To rigorously evaluate personalized agents, we must move beyond conventional task-centric metrics and adopt an evaluation framework that reflects the inherently user-contingent nature of personalization. Unlike standard AI systems, where performance is primarily measured by objective task success, personalized agents are optimized for subjective utility: their behavior should align with an individual user’s preferences, remain consistent across contexts and over time, and adapt as preferences are revealed or revised through interaction, while operating within safety and privacy constraints. Motivated by these considerations, we organize evaluation metrics for personalized agents into five complementary dimensions: Effectiveness, Adaptivity, Generalization, Robustness, and Risk, as illustrated in Table 1. Together, these dimensions provide a structured and holistic view of personalization quality, enabling systematic comparison across diverse tasks, preference sources, and evaluation protocols.
Goal Metric Description
Discovery Accuracy (Li et al., 2025g; Jiang et al., 2024a) Measures whether the agent can correctly discover user preferences or intents from the interaction.
Knowledge Integration Score (Au et al., 2025; Wang et al., 2023a) Measures the ability of the model to incorporate user-specific knowledge or persona information into generated outputs.
Preference Alignment (Zhao et al., 2025a; Wu et al., 2025e) Measures the alignment between the agent’s output and user preferences or stated constraints.
Effectiveness Preference-Aware Planning Accuracy (Xu et al., 2025b, c) Measures whether multi-step plans or tool-use sequences conform to individual user preferences.
Preference Lift (Li et al., 2025g; Hao et al., 2025; Wu et al., 2024c) Compares personalized outputs against a non-personalized baseline on the same prompt, reporting the lift in preference score.
Consistency Score (Mukhopadhyay et al., 2025) Measures the consistency of the agent’s textual style and expressed personality across the whole conversation.
Emotional Empathy (Wu et al., 2025e) Measures the degree to which the response shows understanding and compassion for the user’s emotional state.
Adaptation Success Rate (Shang et al., 2025; Jiang et al., 2025b) Assesses the ratio of successful personalization after receiving new preference signals.
Selective Forgetting (Hu et al., 2025) Measures the ability to revise, overwrite, or remove previously stored information of users when faced with contradictory evidence.
Adaptivity Interaction Efficiency (Li et al., 2025g) Measures the efficiency of adapting, such as number of turns, clarification questions, or edits needed to satisfy user expectations.
Proactivity (Hao et al., 2025) Measures the ability of the agent to proactively identifying user needs and offering extra suggestions or actions to enhance user satisfaction.
Generalization Out-of-Domain Performance (Jiang et al., 2025b) Measures the ability of the agent to generalize user preference to other scenarios or tasks without explicit instruction.
Robustness Accuracy under Ambiguity (Huang et al., 2025d; Pakhomov et al., 2025) Measures the accuracy of the agent’s action when some information is missing or ambiguous.
Misinformed Condition Performance (Huang et al., 2025a) Evaluates the model’s ability to identify and reject incorrect information when the user is misinformed.
Safety Rate (Wu et al., 2025b) Measures whether the agent can make safe decisions based on different user profiles.
Risk Risk Sensitivity (Wu et al., 2025e) Measures whether the agent can recognize and respond appropriately to potential risks in the user’s context.
Privacy Leakage Rate (Mukhopadhyay et al., 2025; Zharmagambetov et al., 2025) Measures the percentage of conversations where the agent makes disclosure in the tasks related to users’ privacy information.
Over-Secrecy Rate (Mukhopadhyay et al., 2025) Measures the failures of utility that arise from excessive caution.
Table 1. Evaluation metrics for personalized agents.
Effectiveness
Effectiveness is the primary criterion for personalization quality and captures whether an agent can deliver user-contingent utility rather than generic helpfulness. It includes (i) preference discovery, where the agent must infer what the user wants from partial interaction traces or contextual signals (Discovery Accuracy (Jiang et al., 2024a)), (ii) preference usage, where agents should correctly incorporates user-specific information into generation (Knowledge Integration Score (Au et al., 2025; Wang et al., 2023a)), and (iii) preference satisfaction, where the final response should comply with explicit constraints and implicit preferences (Preference Alignment (Zhao et al., 2025a; Wu et al., 2025e), Preference-Aware Planning Accuracy (Xu et al., 2025b, c)). Beyond absolute alignment, Preference Lift (Li et al., 2025g; Hao et al., 2025; Wu et al., 2024c) quantifies gains over a non-personalized baseline under the same prompt. In multi-turn settings, Consistency Score (Mukhopadhyay et al., 2025) evaluates stability of persona and values over time, while Emotional Empathy (Wu et al., 2025e) captures affect-appropriate supportive responses.
Adaptivity
Adaptivity evaluates whether a personalized agent can update its behavior as new evidence about preferences arrives while preserving continuity and minimizing user burden. Adaptation Success Rate (Shang et al., 2025; Jiang et al., 2025b) measures whether the agent improves personalization after receiving explicit feedback, new constraints, or additional user context. Selective Forgetting (Hu et al., 2025) that measures the ability to overwrite or retract outdated memory entries when faced with contradictory evidence is equally important due to the evolution of user preferences. Interaction Efficiency (Li et al., 2025g) captures the cost of adaptation, such as the number of turns, clarifications, or edits needed before the user accepts the result, which directly impacts practical usability. Proactivity (Hao et al., 2025) complements these metrics by measuring whether the agent can anticipate needs and offer helpful next steps at the right time without introducing irrelevant suggestions or unnecessary interruptions.
Generalization
Generalization captures whether personalization transfers beyond the narrow contexts where it was learned. Out-of-Domain Performance (Jiang et al., 2025b) measures whether an agent can apply a user’s preferences in specific scenarios to new domains, tasks, or situations without requiring the user to restate them. This is particularly important for agents that aim to support diverse daily activities, where preferences learned in one domain should carry over to other domains in a principled way.
Robustness
Personalized agents operate under inherently noisy and incomplete preference evidence, so a useful evaluation must measure reliability under distributional stress rather than only average-case alignment. Two robustness pressures are especially consequential. Firstly, user preferences are frequently underspecified, partially observed, or expressed indirectly. Therefore, Accuracy under Ambiguity (Huang et al., 2025d; Pakhomov et al., 2025) evaluates whether the agent can make reasonable inferences, ask targeted clarification questions, and avoid overconfident assumptions when key signals are missing. Secondly, user inputs can be wrong or based on misconceptions, either unintentionally or due to stale information. Misinformed Condition Performance (Huang et al., 2025a) captures whether the agent can identify incorrect premises, provide corrective guidance, and still preserve a personalized interaction style while avoiding mistaken beliefs.
Risk
Risk metrics quantify safety, privacy, and other downside constraints that are inseparable from personalization. Key metrics include: Safety Rate (Wu et al., 2025b) measures whether the agent remains safe across heterogeneous user profiles, including cases where preferences might conflict with policies or where the user requests risky actions. Risk Sensitivity (Wu et al., 2025e) evaluates whether the agent recognizes contextual hazards and calibrates its advice, for example by providing warnings, suggesting safer alternatives, or encouraging professional help when appropriate. Because personalization relies on user-specific information, Privacy Leakage Rate (Mukhopadhyay et al., 2025; Zharmagambetov et al., 2025) measures whether the agent discloses private data in disallowed ways, especially under adversarial or indirect prompts. Conversely, Over-Secrecy Rate (Mukhopadhyay et al., 2025) measures the utility loss from excessive caution, where the agent refuses or withholds benign, helpful information despite the user’s legitimate request. Together, these metrics target the core trade-off in personalized agents that aims to maximize user utility subject to risk constraints.
7.2. Assessment Paradigms
The above metrics are instantiated through different assessment paradigms depending on whether the target criterion is objectively verifiable or user-contingent. We group mainstream paradigms into four families. (1) Automatic scoring applies when ground-truth labels or reference answers exist, for example accuracy, precision/recall/F1 scores for classification, and overlap-based metrics such as BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004) for generation. (2) Rule-based constraint checking evaluates compliance with verifiable rubrics or hard constraints, such as whether required attributes are included, forbidden attributes are avoided, or preference constraints are satisfied in structured outputs. This protocol is particularly useful for delegation-style tasks where constraints can be deterministically validated. (3) Learned LLM-based evaluators (LLM-E) fine-tune evaluation models to provide more stable, fine-grained diagnostic feedback for specific preference dimensions, trading off generality for higher consistency in the target domain (Wang et al., 2024a). (4) LLM-as-a-judge (LLM-J) is commonly used for user-contingent criteria such as preference alignment, tone, and satisfaction, where the judge model is conditioned on a user profile or preference description and asked to rate or rank candidate outputs (Zheng et al., 2023). To improve reliability, LLM-J evaluations often adopt pairwise comparisons, calibrated rubrics, and multi-judge aggregation.
7.3. Benchmark
Building on the above metric taxonomy, we summarize representative benchmarks for personalized agents and organize them into two broad families. Interactive Alignment Benchmarks evaluate an agent’s ability to elicit, negotiate, and refine user preferences through multi-turn interaction, where the user remains an indispensable part of the loop. In contrast, User-Substitution Benchmarks evaluate whether an agent can stand in for the user by simulating their preferences or persona to produce user-consistent responses or decisions without further user input. Mainstream benchmarks are summarized in Table LABEL:tab:benchmark-summary [^1] Some benchmarks can partially span both families, and we place them in the major category to avoid duplication..
7.3.1. Interactive Alignment Benchmarks
Preference Discovery Benchmarks
A core challenge in interactive alignment is preference discovery where the agent must recognize, infer, and iteratively refine user intents and preferences from heterogeneous signals, ranging from explicit instructions to implicit feedback and contextual cues (Tsaknakis et al., 2025). In this category, IndieValueCatalog (Jiang et al., 2024a) curates value-expressing statements from diverse individuals and evaluates whether models can infer individualized value judgments, highlighting limitations in capturing fine-grained personal values. PersonaBench (Tan et al., 2025a) constructs synthetic yet realistic personal profiles and associated artifacts to test whether agents can extract and reason over user-specific information. PrefDisco (Li et al., 2025g) proposes a meta-evaluation framework that emphasizes interactive elicitation efficiency, measuring whether an agent can proactively ask informative questions for a user-aligned solution with minimal interaction turns.
Output-level Alignment Benchmarks
This category evaluates personalization in an end-to-end manner, focusing on whether the agent’s final outputs satisfy a target user’s preferences and constraints in realistic interactive settings. For conversation-centered personalization, benchmarks such as ALOE (Wu et al., 2024c) and PrefEval (Zhao et al., 2025a) test whether agents can infer, retain, and consistently follow user preferences across multi-turn dialogues. Beyond conversational tasks, PDR-Bench (Liang et al., 2025) evaluates whether agents can incorporate user profiles into research workflows and produce outputs that are practically useful for the end user. For recommendation, AgentRecBench (Shang et al., 2025) and RecBench+ (Huang et al., 2025a) assess whether agents can generate preference-conditioned recommendations and adapt to user feedback. More generally, PersonaLens (Zhao et al., 2025d) targets multi-scenario, task-oriented assistants and evaluates whether personalization remains coherent across diverse domains and user intents.
Component-probing Benchmarks
Complementary to the above, component-probing benchmarks provide controlled testbeds that isolate specific personalization capabilities, enabling diagnostic analysis of when and how an agent leverages user information. PersonaFeedback (Tao et al., 2025) pairs predefined personas with queries and asks the model to select persona-consistent responses, offering a direct probe of preference-conditioned alignment. A major sub-line focuses on personalized memory, including LongMemEval (Wu et al., ), LoCoMo (Maharana et al., 2024), PerLTQA (Du et al., 2024), MemoryAgentBench (Hu et al., 2025), ConvoMem (Pakhomov et al., 2025), and PAL-Bench (Huang et al., 2025f), which test whether agents can organize, retrieve, and apply user-specific memories under long conversations or incremental preference revelation. For personalized tool use, ETAPP (Hao et al., 2025) and PTBench (Huang et al., 2025d) evaluate whether agents tailor tool-invocation strategies to user needs, including proactivity and handling missing or ambiguous information. Some benchmarks explicitly target adaptivity over time; for example, PersonaMem and PersonaMem-v2 (Jiang et al., 2025a, b) assess whether an agent can track evolving traits and preferences, revise outdated beliefs, and generalize them to new scenarios. Finally, several benchmarks diagnose user-specific risk, such as PSG-Agent (Wu et al., 2025b) and PENGUIN (Wu et al., 2025e) that evaluate profile-contingent safety considerations for LLM-based agents, while PrivacyBench (Mukhopadhyay et al., 2025) and AgentDam (Zharmagambetov et al., 2025) test privacy compliance for agents when handling sensitive user information.
7.3.2. User-Substitution Benchmarks
Textual and Dialogue Benchmarks
Textual and dialogue personalization benchmarks focus on evaluating how models generate content that aligns with individual user preferences, styles, and contextual knowledge. For instance, long-form generation tasks, as in LaMP (Salemi et al., 2023) and LongLaMP (Kumar et al., 2024), test whether models can adapt outputs to evolving user interests over time, while PEFT-U (Clarke et al., 2024) reformulates classification problems into annotator-specific instances to measure user-conditioned generation capabilities. Benchmarks that leverage retrieval or knowledge graphs, such as PGraphRAG (Au et al., 2025), further examine performance in sparse or cold-start scenarios. Beyond single-turn generation, multi-turn conversational benchmarks, including PersonaConvBench (Li et al., 2025e), PER-CHAT (Wu et al., 2021), LaMP-QA (Salemi and Zamani, 2025b), DPL (Qiu et al., 2025b), REGEN (Sayana et al., 2025), and KBP (Wang et al., 2023a)—assess whether models maintain coherence, adapt to a user’s conversational style, and integrate personalized knowledge consistently across dialogue. PRISM (Kirk et al., 2024) targets more complex scenarios where user values vary culturally or contextually, evaluating the reproduction of individualized subjective choices.
Planning and Tool-Use Benchmarks
Benchmarks in planning and tool-use examine whether personalized agents can learn and apply user preferences to perform sequential or multi-step tasks. Preference-based Planning (PBP) (Xu et al., 2025b) simulates everyday activities across diverse environments to evaluate preference-driven planning, while PEToolBench (Xu et al., 2025c) specifically measures the agent’s ability to select and use tools according to individual user preferences. FamilyTool (Wang et al., 2025g) extends this focus to complex tool-based workflows, emphasizing the integration of user-specific needs. Broader multi-domain planning benchmarks, including TravelPlanner++ (Singh et al., 2024), TripTailor (Wang et al., 2025a), TripCraft (Chaudhuri et al., 2025), COMPASS (Qin et al., 2025), TripTide (Karmakar et al., 2025), and Personal Travel Solver (PTS) (Shao et al., 2025), assess tasks ranging from adaptive itinerary generation to disruption handling, testing whether agents can interpret explicit or inferred preferences and construct coherent, user-aligned plans.
Embodied and Spatial Benchmarks
Embodied and spatial benchmarks investigate whether agents can navigate, manipulate objects, and ground actions in physical or simulated spaces according to user preferences. Memory-guided interaction tasks, such as MEMENTO (Kwon et al., 2025), assess the agent’s ability to recall user-specific cues, while PersONAL (Ziliotto et al., 2025) focuses on personalized navigation and object grounding in photorealistic home environments. Personalized Instance-based Navigation (PIN) benchmark (Barsellotti et al., 2024) adds further complexity by requiring agents to locate user-specific target objects among distractors in 3D scenes, emphasizing individualized navigation strategies.
Table 2. Summary of Personalized Benchmark.
Benchmark Scale Task Pref. Goal Eval. Metrics
Interactive Alignment Benchmarks
IndieValueCatalog (Jiang et al., 2024a) 800 Human Value Inferred Effectiveness Auto. Accuracy
PersonaBench (Tan et al., 2025a) 582 General Inferred Effectiveness Auto. Recall, F1
PrefDisco (Li et al., 2025g) 10,000 General Interactive Effectiveness, Adaptivity Auto., Rule-based Discovery Accuracy, Preference Alignment, Interaction Efficiency, Correctness
PersonalLLM (Wu et al., 2024c) 1000 General Given, Inferred Effectiveness LLM-J, LLM-E Personalized Reward
ALOE (Wu et al., 2024c) 100 General Inferred Effectiveness LLM-J Alignment Level, Improvement Rate
PrefEval (Zhao et al., 2025a) 3,000 General Given, Inferred Effectiveness Auto., Rule-based, LLM-J Accuracy, Personalized Rate
AgentRecBench (Shang et al., 2025) 1,500 Recomm. Inferred Effectiveness, Adaptivity Auto. Hit Rate@N
RecBench+ (Huang et al., 2025a) 34,494 Recomm. Given, Inferred Effectiveness, Robustness Auto. Precision, Recall, Condition Match Rate, Fail to Recommend
PDR-Bench(Liang et al., 2025) 250 Deep Research Given Effectiveness LLM-J Personalization Alignment, Content Quality, Factual Reliability
PersonaLens (Zhao et al., 2025d) 122,133 Multi-domain Inferred Effectiveness LLM-J Task Completion Rate, Personalization, Naturalness, Coherence
PersonaFeedback (Tao et al., 2025) 8,298 General Given Effectiveness Auto. Accuracy
PerLTQA (Du et al., 2024) 8,593 Memory Given, Inferred Effectiveness Auto., LLM-J Accuracy, Precision, Recall, F1, Recall@K, MAP, Coherence
LoCoMo (Maharana et al., 2024) 7,512 Memory Inferred Effectiveness, Robustness Auto. Precision, F1, Recall@K, ROUGE
LongMemEval (Wu et al., ) 500 Memory Inferred Effectiveness Auto., LLM-J Accuracy, Precision, Recall, F1, Recall@K, MAP, Coherence
MemoryAgentBench (Hu et al., 2025) 146 Memory Inferred Effectiveness, Adaptivity Auto. Accuracy, Recall@5, F1 in Accurate Retrieval, Test-Time Learning, Long-range Understanding and Selective Forgetting settings
ConvoMem (Pakhomov et al., 2025) 75,336 Memory Inferred Effectiveness, Adaptivity Auto. Accuracy, Cost, Latency
PAL-Bench (Huang et al., 2025f) 100 Memory Inferred Effectiveness Auto., LLM-J BLEU, Win Rate
PTBench (Huang et al., 2025d) 1,083 Tool Usage Given, Inferred Effectiveness, Robustness Auto. Accuracy
ETAPP (Hao et al., 2025) 800 Tool Usage Inferred Effectiveness, Adaptivity Rule-based, LLM-J Personalization Score, Proactivity Score
PersonaMem (Jiang et al., 2025a) 5,990 General Inferred Effectiveness, Adaptivity Auto. Accuracy
PersonaMem-v2 (Jiang et al., 2025b) 10,000 General Inferred Effectiveness, Adaptivity Auto. Accuracy
PSG-Agent (Wu et al., 2025b) 2,900 Multi-domain Inferred Risk Auto. Accuracy, Precision, Recall, F1-score
PENGUIN (Wu et al., 2025e) 14,000 Multi-domain Given Risk LLM-J Risk Sensitivity, Emotional Empathy, User-specific Alignment
PrivacyBench (Mukhopadhyay et al., 2025) 478 General Given Risk LLM-J Leakage Rate, Over-Secrecy Rate, Inappropriate Retrieval Rate, Consistency Score
User-Substitution Benchmarks
LaMP (Salemi et al., 2023) 25,095 General Inferred Effectiveness Auto. Accuracy, F1, MAE, RMSE, ROUGE
LongLaMP (Kumar et al., 2024) 9,658 General Inferred Effectiveness Auto. ROUGE, METEOR
PEFT-U (Clarke et al., 2024) 15,300 General Inferred Effectiveness Auto. BLEU, ROUGE-L, Distinct-1/2
PERSONA (Castricato et al., 2025) 3,868 Dialogue Inferred Effectiveness Auto., LLM-J Perplexity, BLEU, BERTScore, Persona consistency
PGraphRAG (Au et al., 2025) 10,000 Text Gen. Inferred Effectiveness Auto. MAE, RMSE, ROUGE, METEOR
PersonaConvBench (Li et al., 2025e) 111,634 Multi-domain Inferred Effectiveness Auto., LLM-J Accuracy, ROUGE, METEOR, BLEU, SBERT similarity
PER‑CHAT (Wu et al., 2021) 1,500,000 General Inferred Effectiveness Auto. Perplexity, BLEU
LaMP‑QA (Salemi and Zamani, 2025b) 29,666 Q&A Given Effectiveness Auto. ROUGE‑1, ROUGE‑L, METEOR
DPL (Qiu et al., 2025b) 9,472 Text Gen Inferred Effectiveness Auto. ROUGE-1, ROUGE-L, BLEU, METEOR
REGEN (Sayana et al., 2025) 1,258,224 Multi-domain Inferred Effectiveness, Generalization Auto. Recall@K, NDCG@K, MRR, BLEU, ROUGE-L, Semantic similarity
KBP (Wang et al., 2023a) 9,821 Q&A Given Effectiveness Auto., LLM-J BLEU, ROUGE, Persona grounding, Knowledge grounding
PRISM (Kirk et al., 2024) 1,500 Decision Given Effectiveness LLM-J Preference agreement(LLM-J)
PBP (Xu et al., 2025b) 50,000 Planning Inferred Effectiveness Auto. Levenshtein distance
PEToolBench (Xu et al., 2025c) 3,000 Tool Usage Inferred Effectiveness, Adaptivity Auto. Tool Accuracy
TravelPlanner++ (Singh et al., 2024) 1,000 Travel Planning Inferred Effectiveness Auto. Feasibility, Constraint satisfaction
TripTailor (Wang et al., 2025a) 3,800 Travel Planning Inferred Effectiveness, Adaptivity Rule-based, Auto., LLM-J Feasibility, Rationality, Personalization
TripCraft (Chaudhuri et al., 2025) 1,000 Travel Planning Inferred Effectiveness, Adaptivity Auto., LLM-J Temporal, Spatial, Ordering, Persona
TripTide (Karmakar et al., 2025) 1,000 Travel Planning Inferred Effectiveness, Adaptivity Auto., LLM-J Intent preservation, Adaptability
PTS (Shao et al., 2025) 1,000 Travel Planning Inferred Effectiveness, Adaptivity Auto. Constraint satisfaction, Preference alignment
PersonalWAB (Cai et al., 2025) 38,000 Web Search Inferred Effectiveness Auto. Search accuracy, Rec. accuracy
PersONAL (Ziliotto et al., 2025) 1,800 Embodied Nav Inferred Effectiveness, Adaptivity Auto., LLM-J Success, SPL, Grounding
PIN (Barsellotti et al., 2024) 1,193 Embodied Nav Inferred Effectiveness, Adaptivity Auto., LLM-J Success, Path efficiency
MEMENTO (Kwon et al., 2025) 1,900 Embodied Tasks Inferred Effectiveness, Adaptivity Auto., LLM-J Memory recall, Task success
FamilyTool (Wang et al., 2025g) 1,152 Tool Usage Inferred Effectiveness, Adaptivity Auto. Tool accuracy
AgentDam (Zharmagambetov et al., 2025) 246 Web Search Given Risk LLM-J Privacy Leakage Rate
8. Applications
8.1. Conversational Assistant
8.1.1. Daily Dialogue Assistants
Personalized dialogue assistants primarily focus on long-term conversational continuity, where the agent must preserve user-specific context and generate preference-consistent responses across multiple interactions. These agents serve a wide range of functions, from general-purpose conversational support to open-domain recommendation and information retrieval (Zhang et al., 2025h; Huang et al., 2025a). Recent research has begun to treat long-term interaction as a benchmarked personalization problem (Pakhomov et al., 2025; Shang et al., 2025). Jones et al. (2025) further highlights how memory behaviors shape user trust and expectations, emphasizing the need for careful interface design and memory management in deployed personalized assistants. On the product side, chatbots such as ChatGPT (OpenAI, 2022), Gemini (Google, 2023), and DeepSeek (DeepSeek, 2024) have operationalized persistent personalization by allowing users to enable, inspect, and delete stored memories. In recommendation and search systems, responses are further tailored to individual needs based on ongoing user interaction (Kim et al., 2025d; Wang et al., 2024f; Huang et al., 2025e), evolving user-understanding (Wang et al., 2025b; Zhang et al., 2024b), alongside personalized decision-making models (Wang et al., 2024d; Zerhoudi and Granitzer, 2024).
8.1.2. Emotional Support Companions
These agents aim to provide sustained, empathetic assistance for users’ emotional well-being, where effectiveness depends strongly on user-contingent factors such as preferred tone, boundaries, and proactivity (Irfan et al., 2024). Personalization can be realized from both sides, where users may actively shape an agent’s persona and interaction dynamics to better match their needs (Zheng et al., 2025), while agents may adapt their behavior by explicitly modeling supportive strategies during interaction (Alotaibi and Alshahre, 2024). For example, ComPeer (Liu et al., 2024c) provides proactive peer support by detecting key moments in a dialogue and generating supportive check-ins based on conversation history and personalized strategies. MultiAgentESC (Xu et al., 2025e) enables strategy-aware support via a multi-agent pipeline that evaluates dialogue understanding, strategy selection, and response generation, providing a diverse set of personalized responses for complex emotional interactions. ARIEL (Sorino et al., 2024) takes affect-aware support further by integrating physiological signals to track emotional states and adjust responses accordingly. These research trends are reflected in companion-style products such as Replika and Inflection AI’s Pi, which focus on long-term personalization and relationship-building as core product features (Replika, ; AI, ).
8.1.3. Educational Agents
Educational agents use LLM-based systems to enhance both student and educator experiences by providing personalized guidance, adapting instruction to individual progress, and alleviating routine instructional tasks (Sharma et al., 2025; Liang and Tong, 2025). These agents can be divided into two main categories (Chu et al., 2025): (i) Personalized Pedagogical Agents, which automate pedagogical tasks for both teachers and students (Academy, 2025), and (ii) Domain-Specific Agents, which focus on specialized fields like math, music, and language learning, offering tailored support to address domain-specific challenges. Personalized pedagogical agents like EduAgent (Xu et al., 2024), LG-RAG (Cohn et al., 2025), and TeachTune (Jin et al., 2025b) adapt course materials and feedback based on learner profiles and contextual data. Domain-specific agents such as MathAgent (Yan et al., 2025b), Chat Melody (Jin et al., 2025c), Duolingo (Duolingo, 2025), and EduMAS (Li et al., 2024b) integrate domain-specific knowledge to enhance learning outcomes and reduce cognitive load, making learning more efficient and personalized.
8.2. Content Creation
Personalized content creation concerns the generation of standalone textual artifacts, where agents adapt style, structure, and framing to personalization targets rather than interactive dialogue. Typical tasks include personalized writing or summarization with preferences expressed through tone, formality, and rhetorical organization (Xu et al., 2025f; Zhang et al., 2024e; Li et al., 2025h). Personalization is evaluated at the artifact level, giving rise to two complementary paradigms: author-centric alignment with individual writing styles and audience-centric adaptation to reader knowledge and expectations (Salemi et al., 2025; Jangra et al., 2025; Novelo et al., 2025; Cai et al., 2025).
8.2.1. Author-Centric
Author-centric content creation treats personalization as a requirement of authorship rather than an optimization objective, arising in scenarios where users expect generated documents to consistently reflect their individual writing style, habitual phrasing, and subjective preferences. This need is especially prominent in real-world content creation tasks such as creative writing, professional documentation, and brand communication, where users evaluate quality at the artifact level and value stylistic continuity across independent outputs. Accordingly, personalization in this setting is defined by the ability of an agent to act as a stable authorial proxy, producing standalone texts that remain aligned with long-term user identity rather than momentary interaction context (Xu et al., 2025f; Zhang et al., 2024e; Li et al., 2025h). This application perspective is reflected in deployed systems such as Adobe Firefly, which adapts creative content generation to a user’s assets and historical editing patterns (3), Jasper Brand Voice, which operationalizes author-centric alignment for consistent brand and personal voice reproduction (68), and Notion AI, which generates documents conditioned on persistent workspace context and prior authorial conventions (146). Existing research further supports this paradigm by demonstrating that stable preference representations and document-level alignment are central to effective personalized content creation beyond conversational settings (Dey et al., 2025; Bu et al., 2025; Maram et al., 2025; Salemi et al., 2025; Song et al., 2026a; Ueda and Takayanagi, 2025).
8.2.2. Audience-Centric
Audience-centric content creation frames personalization around target readers rather than individual authors, focusing on generating textual artifacts that adapt to audience knowledge, expertise, interests, or usage contexts while preserving the intended informational content (Ning et al., 2025b; Duran and Aytekin, 2025; Zhang et al., 2024c). This paradigm is central to real-world applications such as differentiated instructional content for novice versus expert users, culturally adapted documentation, and audience-aware information presentation in web and software systems, where personalization operates at the level of audience groups rather than individual writing styles. In deployed systems, audience-centric personalization underlies large-scale marketing, communication, and outreach platforms, including social media and advertising content generation tailored to audience segments as exemplified by Predis.ai and Omneky Smart Ads (158; 149), as well as personalized sales and marketing communications targeted at specific recipient groups in systems such as SalesForge AI (176). Correspondingly, recent research emphasizes audience alignment, relevance, and comprehension as defining criteria for personalization quality in this setting, reflecting practical demands observed in audience-driven content delivery and assessment scenarios (Shi et al., 2025a; Song et al., 2026b; Tan et al., 2025b; Lyu et al., 2026; Fu et al., 2025; Salvi et al., 2025).
8.3. Delegation Assistants
Delegation assistants are personalized agents that act on behalf of users by internalizing preferences, managing information, planning actions, and supporting complex cognitive workflows. They operate over extended interaction horizons and assume delegated responsibilities aligned with user-specific goals and contexts.
8.3.1. Information Handling
Information handling concerns the long-term management of user-specific knowledge as a delegated cognitive responsibility. Delegation assistants acquire, organize, retain, and surface relevant information across tasks and over extended time horizons, reducing user cognitive load and interaction overhead. Recent work frames information access itself as a personalized service, emphasizing agentic retrieval and memory-centric designs. ARAG (Maragheh et al., 2025) models retrieval as an agent-driven process that dynamically selects and synthesizes information tailored to individual users, while SPARK (Chhetri et al., 2025) extends this paradigm to personalized search through coordinated retrieval and knowledge sharing under user context. Beyond retrieval, effective information handling requires persistent memory maintenance and selective recall. Systems such as Mr.Rec(Huang et al., 2025b) demonstrate how memory and reasoning modules jointly support personalized assistants that retain user-relevant knowledge beyond single tasks, and log-contextualized RAG (Cohn et al., 2025) shows how accumulated interaction histories enable continuity without repeated user input. These directions are reflected in deployed delegation assistants including Rewind AI (Rewind AI, Inc., 2026), Glean (Glean Technologies, Inc., 2026), You.com (You.com, Inc., 2026), Microsoft Copilot (Microsoft Corporation, 2026), and Mem (Mem, Inc., 2026), where personalization emerges through sustained ownership of the user’s informational context, enabling downstream planning and action to operate on a coherent, user-aligned knowledge substrate.
8.3.2. Task Planning
Task-planning delegation assistants aim to translate high-level user goals into executable action sequences while respecting user-specific preferences and contextual constraints. Personalization in this setting is essential for resolving trade-offs and prioritizing subtasks. Representative systems illustrate how personalization manifests in delegated planning. VAIAGE frames personalized travel planning as a collaborative multi-agent process, where specialized agents jointly coordinate subtasks under user-defined requirements (Liu et al., 2025a). Learning-based approaches further demonstrate that explicitly modeling user preferences during planning enables agents to generate plans that better align with individual goals and constraints (Xu et al., 2025b). Beyond single-task settings, collaborative language-agent frameworks have been applied to complex planning scenarios involving multiple constraints, improving flexibility and robustness in delegated workflows (Zhang et al., 2025a). Personalized task planning has also been explored across application domains, including embodied assistants that adapt planning behavior to human preferences in interactive environments (Han et al., 2025a), and social decision-making contexts such as participatory urban planning (Quan and Lee, 2025).
8.3.3. Behavior Adaptation
Behavior adaptation operates over extended time horizons, where agents align their interaction style, initiative, and decision tendencies with a user’s habitual patterns and expectations. It enables reliable delegation by internalizing stable routines while remaining responsive to evolving goals, thereby reducing user supervision and cognitive effort. Representative systems illustrate how such adaptive delegation is realized in practice. MEAgent models persistent user habits in mobile assistant settings, enabling stable delegation across daily activities (Wang et al., 2026), while memory-centric frameworks allow agents to accumulate and reuse user-specific knowledge beyond single interactions (Westhäußer et al., 2025b; Wang et al., 2025f). Test-time personalization approaches such as PersonaAgent show that adaptive behaviors can be realized dynamically by grounding responses in inferred user personas without retraining (Zhang et al., 2025c). As delegation becomes more autonomous, agents increasingly exhibit proactive behaviors, anticipating user needs rather than remaining purely reactive (Sun et al., 2025b). At the same time, empirical studies of LLM-based GUI agents caution that unconstrained behavior adaptation can lead to misalignment with human values, highlighting the need for controllable personalization in real-world delegation assistants (Gebreegziabher et al., 2026).
8.4. Expert Support in Specific Domains
Personalized agents are increasingly deployed as expert assistants in specialized domains such as healthcare, finance, legal services, and research workflows, where user-centric tailoring can improve relevance and usability but must be reconciled with stringent domain constraints and risk-sensitive requirements (Vishwakarma et al., 2025).
8.4.1. Healthcare
In healthcare, personalization is shifting from static question answering to longitudinal, stateful interaction, where effective agents must support a patient trajectory rather than a single-turn diagnosis. Frameworks such as MedChain (Liu et al., 2024a) and MedPlan (Hsu et al., 2025) organize multi-stage clinical pipelines that span history elicitation, symptom clarification, and treatment planning, emphasizing active querying to resolve uncertainty before delivering user-specific guidance. To support continuity of care, many systems maintain persistent user state, which also heightens the need for safeguards against over-accommodation and unsafe agreement (Westhäußer et al., 2025c). In response, constellation-style architectures such as Polaris and ColaCare decouple patient-facing communication from clinical reasoning, pairing an empathetic interface agent with specialist verification modules that check recommendations against guidelines (Mukherjee et al., 2024; Wang et al., 2025h; Jang et al., 2025). For data-intensive workflows, EHRAgent further reduces hallucination risk by grounding personalized EHR analysis in deterministic code execution rather than purely generative reasoning (Shi et al., 2024; Jin et al., 2024). Finally, evaluation environments such as AgentClinic (Schmidgall et al., 2024) highlight the need to stress-test personalized clinical agents with diverse simulated patient behaviors to assess robustness under realistic interaction dynamics (Lyu et al., 2025).
8.4.2. Finance
In financial domains, personalization should balance alignment with individual goals and behaviors against risks arising from cognitive biases and stringent regulatory requirements, making trust calibration and compliance central challenges for personalized agent design. FinPersona-based studies surface a recurring trust–quality paradox, showing that persona-driven personalization can increase user trust even when the underlying advice quality is poorer or riskier (Takayanagi et al., 2025b, a), highlighting that naive mirroring of user preferences may reinforce risky biases. This motivates safety-aware personalization frameworks such as PSG-Agent (Wu et al., 2025b), which incorporate personality-conditioned guardrails to calibrate recommendations. Alongside trust calibration, finance imposes hard compliance constraints that bound what an agent is permitted to recommend or execute. To support controllable reasoning under these constraints, recent systems increasingly adopt modular architectures. FinRobot (Yang et al., 2024) decomposes workflows into specialized roles for market data aggregation, quantitative analysis, and thesis generation, improving traceability of personalized recommendations. In parallel, compliance-centered designs such as GraphCompliance (Chung et al., 2025b) align user context and intended actions with structured regulatory graphs, enabling verifiable reasoning paths that constrain personalization within enforceable policy boundaries and reduce “plausible” but non-compliant outputs.
8.4.3. Legal services
In legal services, personalization conditions reasoning on user-specific facts, jurisdictions, and drafting intents, transforming legal assistance into a context-sensitive inference problem. A key challenge is statutory drift, where personalized agents must correctly apply evolving laws to individual cases, making static knowledge or one-time verification insufficient. Benchmarks such as LawShift (Han et al., ) show that agents need to adapt updated statutes to the same user-provided facts without retraining, highlighting the need for inference-time adaptability. Personalization also amplifies hallucination risks in legal drafting, as agents tailor arguments and citations to user context. To mitigate this, systems increasingly employ post-hoc verification loops, including Citation-Enhanced Generation (CEG) and watchdog frameworks such as HalMit, which validate personalized drafts against authoritative legal sources before delivery (Li et al., 2024c; Liu et al., 2025b). These are particularly important for multilingual and jurisdiction-aware legal assistants, such as Mina, where personalization must respect local procedural norms while preserving legal correctness (Wasi et al., 2025).
8.4.4. Research Assistance
Research assistance supports expert users across the research lifecycle by performing project- and goal-conditioned scholarly work. Early agentic systems position LLM agents as general research collaborators for exploration and orchestration, exemplified by Agent Laboratory (Schmidgall et al., 2025), while later work moves toward user- and context-conditioned pipelines, including project scoping for specific impact objectives (Emmerson et al., 2025) and adaptive retrieval and dialogue behaviors that reflect individual information-seeking patterns (Mo et al., 2025). Multi-agent settings further develop sustained, stateful workflows through virtual research groups that coordinate iterative investigation (Li et al., 2025a; Naumov et al., 2025). Recent studies emphasize human-centered deployment and interaction design in real research settings (Chen et al., 2025b; Chung et al., 2025a), and specialized assistants instantiated in institutional or scientific domains (Mendoza et al., 2025; Leng et al., 2025) highlight the need to balance autonomy with trust and verifiability in expert use (Zamprogno et al., 2025; Zhang et al., 2025d). Complementary perspectives explore personalized scaffolding and customized writing support for research productivity (Kelley et al., 2025; Kabir et al., 2025; Ahmad, 2025).
9. Open Problems and Future Directions
Despite rapid progress, building robust and deployable personalized agents remains an open challenge. We highlight several key directions that require sustained attention.
9.1. Decision-Critical User Modeling
A fundamental question in personalized agents is how to represent user preferences in a form that is both expressive and actionable. User preferences vary not only across individuals but also across different aspects of a single task, and may even conflict within the same user. Moreover, preferences are rarely stated explicitly and are instead implicitly revealed through historical behavior, interaction patterns, or tool usage. A complementary challenge lies in identifying which aspects of a user are actually decision-critical for a given task, which may differ substantially across application scenarios. General-purpose conversational agents often prioritize broad, high-level alignment with user values and interaction style, whereas domain-specific or high-stakes agents require precise identification of fine-grained attributes that directly affect outcomes. Future work should therefore focus on structured and adaptive preference representations that balance expressiveness with relevance, allowing agents to reason about which user characteristics matter when, rather than assuming a fixed or universally applicable user model.
9.2. Temporal Dynamics and Continual Personalization
Personalization is inherently dynamic and unfolds through sustained interaction rather than being achieved instantaneously. User preferences are progressively revealed as agents observe behavior and feedback, and different aspects of user information evolve at different temporal scales. Short-term interests may shift rapidly across contexts, while long-term values, goals, or reasoning styles tend to remain relatively stable. This temporal heterogeneity poses a fundamental challenge for static personalization mechanisms. Promising directions include continual user modeling techniques that update representations over time without catastrophic forgetting, as well as interaction-aware mechanisms that leverage newly acquired signals to guide subsequent behavior. By closing the feedback loop between observation and action, agents can refine personalization progressively, maintaining alignment as user preferences evolve rather than allowing personalization quality to degrade.
9.3. Generalization
Personalized agents must operate under sparse and heterogeneous user data, raising the question of how personalization can generalize beyond observed interactions. In practice, user-specific signals are often unevenly distributed or unavailable for new users and tasks, making it difficult for agents to reliably align behavior beyond observed interactions. Generalization in personalized agents manifests along two dimensions. First, agents must generalize to unseen or low-interaction users, where limited evidence makes direct preference inference unreliable. Second, agents must generalize across domains and tasks, where preferences learned in one context may only partially apply to another. Even when user intent is stable, differences in task structure, action space, or decision criteria can cause misalignment if personalization mechanisms are overly context-specific. Promising directions include few-shot and meta-learning approaches that rapidly adapt to new users, uncertainty-aware personalization that avoids overconfident alignment when signals are weak, and transferable preference abstractions that capture reusable decision principles rather than task-specific behaviors. Improving generalization is essential for scaling personalized agents beyond narrowly scoped or data-rich scenarios.
9.4. Evaluation Protocols and Benchmarking
Evaluating personalization poses a unique challenge because success is defined relative to individual users rather than objective task completion. Existing benchmarks often rely on synthetic user data generated by language models, which may fail to capture real human diversity and can suffer from homogenization effects. Additionally, common evaluation protocols such as LLM-as-a-judge raise concerns about reliability and alignment with human satisfaction. Future evaluation frameworks should emphasize fine-grained, component-level analysis and distinguish between intra-user consistency and inter-user differentiation. Incorporating human-in-the-loop evaluation, longitudinal interaction studies, and user-centric satisfaction metrics may provide more faithful assessments of personalization quality.
9.5. Privacy and User Control
Personalization inherently relies on collecting and retaining user data, making privacy a central concern for real-world deployment. Even when user data is stored separately, risks remain in the form of unauthorized access, unintended leakage, or misuse of personal information. Moreover, self-evolving agents that reuse historical data for learning may erode user trust if users perceive a loss of control over their data. Future systems must balance data utility with privacy preservation through careful system design. Promising directions include privacy-aware memory architectures, selective retention policies, on-device or federated personalization, and transparent user controls for inspecting and managing stored information. Privacy should be treated not only as a technical constraint but also as a core component of user experience.
9.6. Efficiency and Deployment Cost
A practical question concerns how personalization can be achieved efficiently under real-world constraints. In practice, personalization is often implemented via post-training adaptations or inference-time augmentations applied to a general-purpose agent. These mechanisms introduce additional costs, including memory storage, increased inference latency, and extra computation from re-planning or tool usage. Future work should explore workflow designs that explicitly trade off personalization depth against computational cost. Lightweight personalization mechanisms, adaptive memory usage, and budget-aware inference strategies may enable agents to deliver meaningful personalization without prohibitive overhead.
10. Conclusion
Personalization is a defining capability for next-generation LLM-powered agents, enabling them to move beyond generic task execution toward sustained, user-aligned collaboration. In this survey, we examine personalization from a system-level perspective, organizing existing work around four core capabilities: profile modeling, memory management, planning, and action execution. This decomposition clarifies how user-specific signals are inferred, maintained, and operationalized throughout the agent pipeline, and highlights the inter-dependencies between different components. We further connect these capabilities to real-world deployment by surveying metrics and benchmarks, followed by representative applications. Looking ahead, while substantial progress has been made, personalized agents continue to face fundamental challenges, including sparse and evolving user signals, generalization, evaluation, privacy preservation, and efficiency constraints. Addressing these challenges requires not only advances in individual components, but also holistic agent architectures that balance personalization, robustness, and safety. We hope this survey provides a coherent foundation for understanding the design space of personalized LLM-powered agents, bridges benchmark-driven research with real-world deployment, and inspires future work toward building agents that are capable and reliably aligned with the individuals they serve.